It takes an older pump twice as long to drain a certain pool as it does a newer pump. Working together, it takes the two pumps 4 hours to drain the pool. How long will it take the older pump to drain the pool working alone? Do not do any rounding.
Answer by TimothyLamb(4379)

r = w/t
t = w/r
---
x = rate of newer pump
y = rate of older pump = 1/( 2(1/x) )
y = rate of older pump = 1/( 2/x )
y = rate of older pump = x/2
---
time of newer pump = 1/x
time of older pump = 1/y
---
rate together:
x + y = 1/4
x + x/2 = 1/4
(2/2)x + (1/2)x = 1/4
(3/2)x = 1/4
x = (1/4)/(3/2)
x = (1/4)*(2/3)
x = (1/2)*(1/3)
x = 1/6
y = x/2
y = (1/6)/2
y = 1/12
---
answer:
time of newer pump = 6 hours
time of older pump = 12 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php